### v1

# Quantum-Corrected Thermodynamics of Conformal Weyl Gravity Black Holes: GUP Effects and Phase Transitions

Abstract
We investigate the thermodynamic properties of black holes in Conformal Weyl Gravity (CWG) using the Mannheim-Kazanas solution, with particular emphasis on quantum corrections that become significant near the Planck scale. Our analysis employs the Hamilton-Jacobi tunneling formalism to derive the Hawking temperature, revealing explicit contributions from the conformal parameters $\beta$, $\gamma$, and $k$ that lead to substantial deviations from Schwarzschild black hole behavior. We incorporate quantum gravitational effects through the Generalized Uncertainty Principle, demonstrating systematic suppression of thermal radiation in the near-Planckian regime. Using an exponentially corrected entropy model, we compute the complete spectrum of QC thermodynamic potentials, including internal energy, pressure, heat capacity, and free energies. Our heat capacity analysis reveals critical phase transitions separating thermodynamically stable and unstable regions, with the scale-dependent parameter $\gamma$ playing a crucial role in controlling phase structure. The Joule-Thomson expansion analysis shows distinct cooling and heating regimes with inversion points that shift systematically with CWG parameters, capturing QC phase transitions absent in general relativity. We also examine gravitational redshift in CWG geometry, finding complex radial dependence that could provide observable signatures for distinguishing conformal gravity from Einstein’s theory. Our results demonstrate that CWG offers a consistent framework for studying black hole thermodynamics beyond general relativity, with quantum corrections revealing rich phase structures and potential observational consequences.
Black hole; Joule-Thomson Expansion; Quantum Correction; Redshift; Radiation; Generalized Uncertainty Principle; Conformal Weyl Gravity.
I Introduction
The study of black hole (BH) thermodynamics has emerged as one of the most profound and intellectually rewarding intersections between general relativity (GR), quantum field theory, and statistical physics, fundamentally transforming our understanding of the deep connections between gravity, thermodynamics, and quantum mechanics. Although Einstein’s GR has achieved remarkable success in describing macroscopic gravitational phenomena across vast effective energy scales, from planetary motion to cosmological dynamics, several fundamental conceptual and observational challenges at both infrared and ultraviolet scales have persistently motivated the exploration of alternative theoretical frameworks. These challenges include the notorious cosmological constant problem, the enigmatic galactic rotation curve anomalies that suggest the presence of dark matter, the quantum origin of the BH entropy and the associated information paradox, and the fundamental failure of GR to provide a complete quantum theory of gravity [1, 2].
CWG stands out prominently among alternative gravity theories due to its elegant local conformal invariance and sophisticated higher-derivative mathematical structure [3, 4, 5, 6, 7]. Unlike conventional GR, which constructs its gravitational action from the Ricci scalar $R$, CWG employs the square of the conformal Weyl tensor $C_{\mu\nu\rho\sigma}C^{\mu\nu\rho\sigma}$ as its fundamental building block, thereby introducing fourth-order field equations that admit significantly richer vacuum structures and more complex dynamical behaviors. The Mannheim-Kazanas solution, derived within this theoretical framework, represents a notable departure from classical Schwarzschild geometry through the inclusion of additional linear $\gamma r$ and quadratic $kr^{2}$ potential terms that emerge naturally from higher-order field equations [3, 8]. These non-Einsteinian contributions offer compelling explanations for various astrophysical observations, including galactic rotation curves and large-scale structure formation, without requiring the introduction of exotic dark matter components [9]. Furthermore, the presence of local conformal symmetry in CWE does not tolerate any dimensionful parameter (and hence no rigid scale) at sufficiently high energy scales. So, the Newton constant with dimension $[m^{-2}]$-which causes the irremediable catastrophic infinities at loop-level and hence leads to perturbatively non-renormalizable quantum gravity-comes into existence only in the classical vacuum of CWE where the conformal symmetry is broken.
The thermodynamic properties of BHs within the CWG exhibit fascinating modifications compared to their GR counterparts [10, 11, 12, 13, 14], leading to altered causal structures, modified thermal radiation spectra, and fundamentally different horizon properties that reflect the underlying conformal symmetry [15, 16]. These modifications become particularly pronounced in extreme parameter regimes, where the additional CWG terms dominate over the standard Newtonian contributions, suggesting potential observational signatures that could distinguish CWG from conventional gravity theories through precision astronomical measurements.
The incorporation of quantum corrections into BH thermodynamics represents another crucial frontier in our understanding of gravitational physics near the Planck scale, where classical descriptions inevitably break down and quantum effects become dominant [17, 18]. The Generalized Uncertainty Principle (GUP) [19, 20, 21, 22, 23, 24, 25], which emerges naturally from various approaches to quantum gravity including string theory, loop quantum gravity, and non-commutative geometry, provides a minimal yet powerful framework for capturing leading-order quantum gravitational effects through the introduction of a fundamental minimal length scale [26, 27, 28, 29, 30, 31]. This principle modifies the standard Heisenberg uncertainty relations and leads to systematic corrections in BH thermal properties, particularly affecting temperature, entropy, and heat capacity in ways that become increasingly important as BH masses approach Planck-scale values.
Recent advances in observational astronomy, particularly through gravitational wave detection by LIGO/Virgo collaborations and direct BH imaging by the Event Horizon Telescope, have opened unprecedented opportunities for testing alternative gravity theories and probing quantum gravitational effects in realistic astrophysical contexts [32, 33]. These observational breakthroughs demand corresponding theoretical developments that can provide precise predictions for observable signatures of modified gravity and quantum corrections in BH physics.
The Joule-Thomson expansion (JTE) represents a particularly sensitive probe of BH thermodynamics, providing detailed information about phase transitions and thermal stability properties that are invisible to other thermodynamic measures [34, 35, 36, 37]. In the extended phase space formalism, where the cosmological constant is treated as a thermodynamic pressure, JTE analysis reveals rich phase structures including cooling and heating regimes separated by critical inversion points. These phenomena become even more complex in CWG due to the additional parameter space introduced by the conformal corrections. Gravitational redshift measurements offer another powerful avenue for testing alternative gravity theories, as they provide direct observational access to the spacetime metric structure near massive compact objects [38, 39]. The modification of redshift patterns in CWG compared to GR predictions could potentially be detected through high-precision spectroscopic observations of radiation emanating from the vicinity of supermassive BHs or neutron stars.
Our primary motivation for this comprehensive investigation stems from the recognition that understanding BH physics in alternative gravity theories, particularly when augmented with quantum corrections, represents a crucial step toward developing a complete theory of quantum gravity. CWG provides an ideal theoretical laboratory for this investigation due to its mathematical elegance, observational relevance, and rich phenomenological structure. Using the Mannheim-Kazanas metric as a gravitational background, we systematically investigate the thermodynamic behavior of BHs within CWG, with particular focus on quantum corrections that become significant near the Planck scale. We first revisit the semiclassical Hawking radiation through the Hamilton-Jacobi tunneling method [40, 41] and then incorporate corrections from the GUP, which is widely regarded as a minimal extension capturing leading-order quantum gravity effects [42, 43]. These corrections impact key thermodynamic quantities such as temperature, entropy, and heat capacity, opening possibilities for rich phase structures, including modifications to the JTE and redshift phenomena.
The relevance of this analysis lies in its capacity to offer deeper understanding of BH thermodynamics in a conformally invariant setting, enriched with quantum statistical considerations [44, 45, 46]. Moreover, by exploring the interplay between higher-derivative gravity and minimal-length scenarios, our study contributes toward the broader goal of probing the semi-quantum regime where conventional gravitational and thermodynamic intuitions begin to break down [47, 48]. Such insights could prove instrumental in guiding phenomenological models in quantum gravity and in interpreting upcoming observational data from gravitational-wave astronomy and BH imaging [49, 50, 51]. The specific objectives of our research include the development of a comprehensive thermodynamic framework for CWGBHs that incorporates both the geometric modifications arising from higher-order curvature terms and the quantum corrections emerging from the GUP. We systematically analyze thermal radiation properties, compute QC thermodynamic potentials including energy, entropy, pressure, and heat capacity, and investigate phase transition phenomena through detailed thermodynamic analysis. Additionally, we explore gravitational redshift modifications that could provide observational signatures of CWG effects in astrophysical contexts.
The paper is organized as follows: In Section II, we provide a comprehensive review of the Mannheim-Kazanas BH geometry within the CWG framework, establishing the mathematical foundation for our subsequent analysis. Section III presents our evaluation of Hawking radiation using the Hamilton-Jacobi formalism for tunneling particles, revealing the thermal spectrum modifications introduced by conformal corrections. Section IV is devoted to the systematic calculation of quantum corrections arising from the GUP and their impact on thermal radiation properties. In Section V, we conduct a detailed study of quantum-corrected (QC) thermodynamics for CWGBHs, including the computation of all relevant thermodynamic potentials and their phase structure analysis. Section VI focuses on the QC JTE in CWGBHs, investigating isenthalpic processes and thermal stability properties. Section VII analyzes gravitational redshift phenomena in CWGBH geometry, exploring potential observational signatures of conformal corrections. Finally, Section VIII concludes our investigation by summarizing the main results and outlining prospective research directions that emerge from our findings.
II Review Of CWGBH Spacetime: Analysis of the Mannheim-Kazanas Solution
In CWG, BH solutions emerge from a fundamentally different gravitational action compared to Einstein’s GR. Rather than being based on the Ricci scalar $R$ as in the conventional Einstein-Hilbert formulation, CWG employs the square of the Weyl tensor $C_{\mu\nu\rho\sigma}$ as its foundational building block. This approach leads to a locally conformally invariant action that takes the elegant form [3, 4, 5, 6]
$S=-\alpha_{g}\int d^{4}x\sqrt{-g}\,C_{\mu\nu\rho\sigma}C^{\mu\nu\rho\sigma},$ (1)
where $\alpha_{g}$ represents a dimensionless coupling constant that characterizes the strength of the conformal gravitational interaction. The conformal Weyl tensor in four spacetime dimensions, originally formulated by Weyl and later developed by Bach [52, 53], is explicitly expressed as
$C_{\mu\nu\rho\sigma}=R_{\mu\nu\rho\sigma}-2R_{[\mu[\rho}g_{\sigma]\nu]}+\frac{1}{3}g_{\mu[\rho}g_{\sigma]\nu}R,$ (2)
where $R_{\mu\nu\rho\sigma}$ and $R_{\mu\nu}$ denote the Riemann and Ricci tensors respectively, while $R$ represents the scalar curvature. The conformally invariant Weyl tensor inherits the same fundamental symmetry properties as the ordinary Riemann tensor, satisfying the essential relations
$C_{\mu\nu\rho\sigma}=C_{[\mu\nu][\rho\sigma]}=C_{\rho\sigma\mu\nu},\quad C_{\mu[\nu\rho\sigma]}=0,\quad C^{\mu}{{}_{\nu\mu\sigma}}=0.$ (3)
It is noteworthy that this tensor identically vanishes in 2+12+12 + 1-dimensional spacetimes, where it is replaced by the Cotton tensor, highlighting the special role of four-dimensional geometry in conformal gravity theories [54].
The explicit form of the Weyl square term appearing in the action in Eq.(1) can be decomposed as [^1] By utilizing the topological Gauss-Bonnet invariant in conjunction with Eq.(5), the primary action in Eq.(1) admits an alternative representation: $S=-2\alpha_{g}\int d^{4}x\sqrt{-g}\,(R^{2}_{\mu\nu}-\frac{1}{3}R^{2}).$ (4)
$C_{\mu\nu\rho\sigma}C^{\mu\nu\rho\sigma}=R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-2R_{\mu\nu}R^{\mu\nu}+\frac{1}{3}R^{2}.$ (5)
A crucial distinguishing feature of CWG compared to Einstein-Hilbert gravity lies in its respect for local conformal symmetry. The theory remains invariant under pointwise rescalings of the metric tensor according to $g_{\mu\nu}\rightarrow\Omega^{2}(x)g_{\mu\nu}$, where $\Omega(x)$ represents an arbitrary point-dependent positive function [^2] Under such local conformal transformations, the Weyl tensor transforms as $C_{\mu\nu\rho\sigma}\rightarrow C_{\mu\nu\rho\sigma}^{{}^{\prime}}=\Omega^{2}(x)C_{\mu\nu\rho\sigma}$, preserving the form of the action.. This enhanced symmetry structure has profound implications for the resulting field equations and vacuum solutions [3].
The field equations governing the dynamics of matter fields coupled to CWG take the form
$\frac{1}{\sqrt{-g}}\,\frac{\delta S}{\delta g_{\mu\nu}}=-2\alpha_{g}B^{\mu\nu}=-\frac{1}{2}T^{\mu\nu},$ (6)
where $T_{\mu\nu}$ represents the energy-momentum tensor of matter fields, and $B^{\mu\nu}$ denotes the celebrated Bach tensor, defined through the expression
$B^{\mu\nu}=\nabla_{\alpha}\nabla_{\beta}C^{\mu\alpha\nu\beta}-\frac{1}{2}R^{\alpha\beta}C_{\mu\alpha\nu\beta}.$ (7)
This tensor possesses remarkable mathematical properties: it is symmetric, traceless, and divergence-free, making it the natural generalization of the Einstein tensor to fourth-order gravity theories. Within the CWG framework, Birkhoff’s theorem continues to hold, and the cosmological constant emerges naturally as an integration constant within the vacuum solutions of the model [55].
Due to its gauge-like local symmetry structure, the CWG does not inherently define any characteristic length or mass scale. The spontaneous breaking of local conformal symmetry generates the dimensionful Newton’s constant through a mechanism analogous to the Higgs mechanism in particle physics [56] [^3] Alternative realizations of local Weyl symmetry can be achieved through the introduction of additional abelian gauge and scalar fields within Weyl’s geometric framework. For comprehensive treatments, see [57, 58] for Weyl gauge-invariant quadratic and cubic gravity models, and [59, 60, 61, 62] for the integration of local conformal symmetry with Standard Model gauge symmetries, and [63, 64] (and references therein) for the local Weyl conformal invariant Topologically Massive Gravity.. Recent investigations into the fundamental features and phenomenological consequences of CWG have been extensively documented in the literature [65, 66, 67, 68, 69, 70, 71, 72, 73], revealing its rich mathematical structure and potential astrophysical applications.
A fundamental observation is that in vacuum regions where the Ricci tensor vanishes, the Bach tensor $B^{\mu\nu}$ also vanishes identically. Consequently, all vacuum solutions of Einstein’s field equations automatically satisfy the CWG field equations, though the converse is not generally true, indicating that CWG admits a broader class of vacuum configurations than GR [3] [^4] This can be easily seen as one recast Eq.(7) into the form with Schouten tensor ${\cal P}_{\mu\nu}$ as $B_{\mu\nu}=2\nabla^{\alpha}\nabla_{[\alpha}{\cal P}_{\mu]\nu}-{\cal P}^{\alpha\beta}C_{\mu\alpha\nu\beta},$ (8) with ${\cal P}_{\mu\nu}=(1/2)(R_{\mu\nu}-(1/6)g_{\mu\nu}R)$ [74]. Since ${\cal P}_{\mu\nu}=(1/6)\Lambda g_{\mu\nu}$ in Einstein spaces (with $R_{\mu\nu}=\Lambda g_{\mu\nu}$ and $R=4\Lambda$), all the solutions of Einstein spaces are also those of CWE albeit its invalid reverse. For an informative study, see also [67].
For static, spherically symmetric spacetimes, the most general metric ansatz takes the standard form
$ds^{2}=-A(r)dt^{2}+B(r)dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),$ (9)
where the functions $A(r)$ and $B(r)$ must be determined by solving the appropriate field equations derived from the CWG action.
Through systematic analysis of static, spherically symmetric vacuum solutions to the fourth-order Bach field equations, Mannheim and Kazanas discovered an exact analytical solution that we designate as the CWGBH metric. The temporal and radial metric components are governed by the remarkable function [3, 4, 5, 6]
$A(r)={B(r)}^{-1}=1-3\beta\gamma-\frac{\beta(2-3\beta\gamma)}{r}+\gamma r+kr^{2}.$ (10)
This solution exhibits dramatic departures from the classical Schwarzschild BH geometry of GR. The appearance of both a linear potential term $\gamma r$ and a cosmological term $kr^{2}$ arises directly from the fourth-order nature of the conformal field equations, representing genuine quantum gravitational corrections absent in Einstein’s theory. The parameter $\beta$ functions analogously to the gravitational mass parameter in GR, while $\gamma$ introduces scale-dependent modifications to the gravitational potential, and $k$ behaves as an effective cosmological constant [3].
The integration constant $\gamma$ has received considerable attention in astrophysical contexts, where it has been interpreted as quantifying large-scale deviations from Newtonian gravity. This parameter plays a crucial role in modeling galactic rotation curves without requiring the introduction of dark matter, offering an alternative explanation for observed galactic dynamics. The classical Schwarzschild solution is recovered as a special limiting case when $\gamma\to 0$, $k\to 0$, and $\beta=2GM$.
The radial structure of the metric function reveals distinct physical regimes: the term $-\beta(2-3\beta\gamma)/r$ dominates the Newtonian gravitational potential at short distances, the linear term $\gamma r$ becomes significant at intermediate scales, while the quadratic term $kr^{2}$ governs the behavior at cosmological distances. The zeros of $B(r)$ determine the locations of event horizons, and depending on the relative magnitudes and signs of the parameters $\beta$, $\gamma$, and $k$, the solution can exhibit multiple real, positive roots corresponding to distinct BH horizons.
From a geometric perspective, the Mannheim-Kazanas or CWGBH metric represents a non-trivial vacuum solution of the fourth-order Bach equations, which replace Einstein’s second-order field equations in CWG. These higher-order equations admit more general vacuum configurations due to their enhanced derivative structure and enlarged symmetry content. The associated curvature invariants reveal the presence of curvature singularities at $r=0$, maintaining consistency with classical BH behavior, while the causal structure defined by the horizon locations provides a rich framework for studying thermodynamic properties.
III Hawking Radiation of CWGBHs via Hamilton-Jacobi Formalism of Tunneling Particles
The investigation of quantum tunneling processes for scalar particles in curved spacetime geometries provides a powerful framework for understanding thermal radiation from BHs. The relativistic Hamilton-Jacobi equation offers an elegant and computationally tractable approach to analyze such phenomena within the context of CWG. Given the spacetime geometry characterized by the Mannheim-Kazanas solution in Eq.(10), we examine the behavior of a scalar field described by an action $S$ that satisfies the fundamental constraint [75]
$g^{\mu\nu}\partial_{\mu}S\partial_{\nu}S+m^{2}=0.$ (11)
The inverse components of the metric tensor for the CWGBH spacetime are explicitly given by
$g^{tt}=-\frac{1}{B(r)},\quad g^{rr}=B(r),\quad g^{\theta\theta}=\frac{1}{r^{2}},\quad g^{\phi\phi}=\frac{1}{r^{2}\sin^{2}\theta},$ (12)
where $B(r)$ represents the radial metric function derived from the CWG field equations. Substituting these metric components into the Hamilton-Jacobi equation yields the explicit form
$-\frac{(\partial_{t}S)^{2}}{B(r)}+B(r)(\partial_{r}S)^{2}+\frac{(\partial_{\theta}S)^{2}}{r^{2}}+\frac{(\partial_{\phi}S)^{2}}{r^{2}\sin^{2}\theta}+m^{2}=0.$ (13)
For the analysis of radial tunneling processes, we focus attention on purely radial trajectories of outgoing particles and employ the standard separable ansatz for the action
$S=-Et+W(r)+J_{\theta}\theta+J_{\phi}\phi,$ (14)
where $E$ represents the energy of the tunneling particle, $W(r)$ denotes the radial component of the action, and $J_{\theta}$, $J_{\phi}$ are the angular momentum quantum numbers. For purely radial motion, we set the angular momenta to zero ($J_{\theta}=J_{\phi}=0$), which simplifies the analysis considerably and allows us to focus on the essential physics of the tunneling process.
The temporal and radial derivatives of the action are then given by
$\partial_{t}S=-E,\quad\partial_{r}S=\frac{dW}{dr}.$ (15)
Substituting these expressions into the Hamilton-Jacobi equation and solving for the radial derivative yields
$-\frac{E^{2}}{B(r)}+B(r)\left(\frac{dW}{dr}\right)^{2}+m^{2}=0,$ (16)
which can be rearranged to obtain
$\frac{dW}{dr}=\frac{\sqrt{E^{2}-B(r)m^{2}}}{B(r)}.$ (17)
The radial contribution to the total action is therefore expressed as the integral
[…]
Near the event horizon located at $r=r_{h}$, where the metric function vanishes identically ($B(r_{h})=0$), the integrand develops a singular behavior that requires careful mathematical treatment. To handle this singularity appropriately, we introduce a proper radial coordinate transformation defined by
$d\xi=\frac{dr}{\sqrt{B(r)}}.$ (19)
Expanding the metric function $B(r)$ in a Taylor series near the horizon yields $B(r)\approx B^{\prime}(r_{h})(r-r_{h})$ for small deviations from the horizon radius. Consequently, the proper radial distance behaves as
$\xi\approx\frac{2\sqrt{r-r_{h}}}{\sqrt{B^{\prime}(r_{h})}}.$ (20)
Expressing the action integral in terms of this new coordinate system, we obtain
[…]
For small values of $\xi$ in the near-horizon region, the metric function can be approximated as $B(r(\xi))\approx\frac{B^{\prime}(r_{h})\xi^{2}}{4}$. Focusing on the physically relevant case of massless particles ($m=0$), which dominates the high-energy Hawking radiation spectrum, the integral simplifies dramatically to
$W(\xi)\approx\frac{2E}{\sqrt{B^{\prime}(r_{h})}}\int\frac{d\xi}{\xi}.$ (22)
This integral exhibits a characteristic logarithmic divergence at the horizon, which gives rise to an imaginary contribution to the action. This imaginary part is crucial for the tunneling interpretation and is given by
$\operatorname{Im}W=\frac{2\pi E}{B^{\prime}(r_{h})}.$ (23)
Therefore, the imaginary part of the total classical action becomes
$\operatorname{Im}S=\frac{2\pi E}{B^{\prime}(r_{h})}.$ (24)
According to the semiclassical tunneling framework developed for BH radiation, the emission probability for particles to tunnel through the gravitational potential barrier is governed by [20]
$\Gamma\sim\exp\left(-\frac{4\pi E}{B^{\prime}(r_{h})}\right).$ (25)
This exponential suppression factor directly reveals the thermal nature of the radiation and allows us to identify the Hawking temperature as
$T_{H}=\frac{B^{\prime}(r_{h})}{4\pi}=\frac{3\beta^{2}\gamma}{4\pi r_{h}^{2}}+\frac{\beta}{2\pi r_{h}^{2}}+\frac{\gamma}{4\pi}+\frac{kr_{h}}{2\pi}.$ (26)
This expression represents a significant generalization of the standard Schwarzschild result and demonstrates the rich thermal structure emerging from CWG. The temperature formula is consistent with results obtained through alternative methods such as the null geodesic approach, providing an important cross-validation of the thermal spectrum derivation.
$r_{h}=1$ $r_{h}=5$ $r_{h}=10$
$k$ $\gamma$ $\beta$ $T_{H}$ $\Delta T_{H}$ $k$ $\gamma$ $\beta$ $T_{H}$ $\Delta T_{H}$ $k$ $\gamma$ $\beta$ $T_{H}$ $\Delta T_{H}$
0 0 0.1 0.0159 0 0 0.1 0.00127 0 0 0.1 0.000318
0 0 0.5 0.0796 +400.6% 0 0 0.5 0.00637 +401.6% 0 0 0.5 0.00159 +400.0%
0 0 1.0 0.1592 +901.3% 0 0 1.0 0.0127 +900.0% 0 0 1.0 0.00318 +900.0%
0 1 0.1 0.0963 +505.7% 0 1 0.1 0.0817 +6334% 0 1 0.1 0.0806 +25257%
0 1 0.5 0.1599 +905.7% 0 1 0.5 0.0868 +6733% 0 1 0.5 0.0819 +25677%
0 1 1.0 0.2396 +1406% 0 1 1.0 0.0932 +7240% 0 1 1.0 0.0836 +26195%
1 0 0.1 0.1759 +1006% 1 0 0.1 0.08127 +6298% 1 0 0.1 0.1593 +49906%
1 0 0.5 0.2396 +1406% 1 0 0.5 0.08637 +6699% 1 0 0.5 0.1606 +50315%
1 0 1.0 0.3192 +1907% 1 0 1.0 0.0927 +7200% 1 0 1.0 0.1623 +50849%
1 1 0.1 0.2563 +1512% 1 1 0.1 0.1627 +12717% 1 1 0.1 0.2399 +75377%
1 1 0.5 0.3199 +1912% 1 1 0.5 0.1678 +13118% 1 1 0.5 0.2412 +75786%
1 1 1.0 0.3996 +2412% 1 1 1.0 0.1742 +13622% 1 1 1.0 0.2429 +76321%
Table 1: Hawking temperature $T_{H}$ for CWGBHs across various parameter combinations of horizon radius $r_{h}$, cosmological parameter $k$, scale-dependent correction $\gamma$, and mass-like parameter $\beta$. The $\Delta T_{H}$ column shows percentage increases relative to the baseline case ($k=0$, $\gamma=0$, $\beta=0.1$) for each horizon radius, demonstrating the significant thermal enhancement effects introduced by CWG parameters.
Table 1 presents a comprehensive analysis of the Hawking temperature $T_{H}$ across different parameter regimes, revealing the complex interplay between the horizon radius $r_{h}$, the effective mass-like parameter $\beta$, the scale-dependent correction $\gamma$, and the cosmological analog $k$. The data demonstrates several key physical insights: For compact BHs with small horizon radii ($r_{h}=1$), the inverse square terms in Eq.(26) dominate the thermal behavior, making $T_{H}$ extremely sensitive to variations in both $\beta$ and $\gamma$ parameters. The percentage increases shown in the $\Delta T_{H}$ column illustrate that even modest changes in these parameters can lead to dramatic thermal enhancements, with temperature increases exceeding 2000% in some parameter combinations.
As the horizon radius increases to intermediate values ($r_{h}=5$) and large scales ($r_{h}=10$), the relative importance of the linear $kr_{h}$ term becomes increasingly pronounced, particularly when $k=1$. This transition reflects the changing balance between local gravitational effects (governed by $\beta$ and $\gamma$) and global spacetime geometry (controlled by $k$). The parameter $\beta$ consistently contributes to temperature enhancement across all scales, confirming its interpretation as a gravitational mass analog, while $\gamma$ introduces both direct constant contributions and nonlinear coupling effects with $\beta$, emphasizing its role in modifying the characteristic length scales of the theory.
IV GUP Corrected Hawking Radiation of CWGBHs
To systematically account for the profound influence of quantum gravitational effects on BH thermodynamics, one of the most promising and well-developed theoretical approaches involves the implementation of the GUP. This principle introduces a fundamental minimal length scale that emerges naturally from several candidate theories of quantum gravity, including string theory, loop quantum gravity, and non-commutative geometry [76, 77, 78]. The GUP framework provides a phenomenological window into the quantum gravitational regime by modifying the standard Heisenberg uncertainty relations through the incorporation of Planck-scale corrections.
The modified uncertainty relation in its most commonly employed form is expressed as [79, 80, 81]
$\Delta x\Delta p\geq\frac{\hbar}{2}\left(1+\lambda^{2}l_{p}^{2}\frac{(\Delta p)^{2}}{\hbar^{2}}\right),$ (27)
where $\lambda$ represents a model-dependent dimensionless constant typically of order unity, and $l_{p}=\sqrt{\hbar G/c^{3}}$ denotes the fundamental Planck length scale. This deformation of the uncertainty principle can be inverted to express the momentum uncertainty in terms of the position uncertainty, yielding
$\Delta p\geq\frac{\hbar}{2\Delta x}\left(1+\frac{2\lambda^{2}l_{p}^{2}}{\Delta x^{2}}\right).$ (28)
In the context of BH physics, particularly near the event horizon where quantum effects become most pronounced, it is physically reasonable to associate the fundamental position uncertainty with the characteristic horizon length scale, leading to the identification $\Delta x\sim 2r_{h}$. This association reflects the fact that the horizon represents the natural boundary beyond which classical notions of spacetime geometry begin to break down [82, 83].
The energy of particles emitted through quantum tunneling processes, as estimated from the uncertainty principle applied to momentum fluctuations, receives quantum gravitational corrections according to
$E_{\text{GUP}}\sim\Delta p\sim\frac{\hbar c}{4r_{h}}\left(1+\frac{\lambda^{2}l_{p}^{2}}{2r_{h}^{2}}\right).$ (29)
This fundamental modification in the effective energy of tunneling particles directly impacts the semiclassical tunneling probability that governs Hawking radiation emission. In the standard semiclassical framework, the emission probability follows the characteristic exponential suppression [27]
$\Gamma\sim\exp\left(-\frac{2\pi E}{\kappa}\right),$ (30)
where $\kappa=\frac{1}{2}B^{\prime}(r_{h})$ represents the surface gravity evaluated at the horizon. Under the influence of GUP corrections, this tunneling probability becomes modified to
$\Gamma_{\text{GUP}}\sim\exp\left(-\frac{2\pi E_{\text{GUP}}}{\kappa}\right)=\exp\left(-\frac{4\pi E}{B^{\prime}(r_{h})}\left(1+\frac{\lambda^{2}l_{p}^{2}}{2r_{h}^{2}}\right)\right).$ (31)
From this modified exponential structure, we can directly extract the QC Hawking temperature by identifying the effective thermal factor in the Boltzmann-like distribution. This yields the GUP-corrected temperature as
$T_{\text{GUP}}=\frac{B^{\prime}(r_{h})}{4\pi}\left(1+\frac{\lambda^{2}l_{p}^{2}}{2r_{h}^{2}}\right)^{-1}=T_{H}\left(1+\frac{\lambda^{2}l_{p}^{2}}{2r_{h}^{2}}\right)^{-1},$ (32)
where $T_{H}$ represents the classical Hawking temperature derived in the previous section. This result demonstrates that the emission spectrum undergoes a systematic redshift due to quantum gravitational corrections, with the magnitude of the effect being inversely related to the square of the horizon radius.
The physical interpretation of this correction is particularly illuminating: as the BH horizon radius approaches the Planck scale ($r_{h}\to l_{p}$), the quantum gravitational corrections become increasingly pronounced, effectively suppressing the radiation temperature compared to its semiclassical counterpart. While these effects remain negligible for macroscopic astrophysical BHs with horizon radii many orders of magnitude larger than the Planck length, they become critically important when investigating microscopic BHs, primordial BH evolution, or the final stages of BH evaporation where the horizon approaches Planck-scale dimensions. The GUP framework thus provides an essential conceptual and computational bridge between standard semiclassical thermodynamic descriptions and the quantum gravitational regimes where conventional field theory breaks down.
V QC Thermodynamics of CWGBHs
In the regime where BH dimensions approach the Planck scale, classical thermodynamic descriptions become fundamentally inadequate, necessitating the incorporation of quantum corrections that capture the underlying microscopic degrees of freedom. One of the most sensitive thermodynamic quantities to such quantum effects is the entropy, which in the classical Bekenstein-Hawking framework is simply proportional to the event horizon area. However, this classical description fails to account for the rich quantum structure that emerges near the Planck scale [84].
To provide a more complete theoretical framework, quantum statistical mechanics offers a sophisticated approach through systematic counting of horizon microstates [44, 45]. In canonical ensembles with fixed energy and particle number, quantum fluctuations give rise to subleading modifications to the classical entropy expressions. These corrections effectively capture the statistical uncertainties arising from quantum degrees of freedom that are entirely absent in semiclassical treatments, providing crucial insights into the quantum nature of gravitational thermodynamics [85, 86].
One particularly well-motivated proposal involves the exponentially corrected entropy form, which naturally emerges from specific microstate counting schemes in various approaches to quantum gravity [46, 87].
$S=S_{0}+e^{-S_{0}},$ (33)
where $S_{0}$ represents the classical Bekenstein-Hawking entropy. The exponential correction term becomes increasingly relevant for small values of $S_{0}$, which corresponds precisely to near-extremal or Planck-scale BHs where quantum effects are expected to dominate.
Using this entropy modification within the framework of CWG, the corrected internal energy can be systematically computed by applying the first law of BH thermodynamics
[…]
where $T_{H}$ denotes the Hawking temperature derived in Eq.(26). Substituting the QC entropy expression and performing the integration yields
$E\approx\frac{1}{5}\pi k\,r_{h}^{5}+\frac{1}{8}\pi\gamma r_{h}^{4}-\frac{3}{4}\pi\,\beta^{2}\gamma r_{h}^{2}+\frac{1}{2}\pi\beta r_{h}^{2}.$ (35)
Figure 1 presents a comprehensive density plot illustrating the behavior of the internal energy $E$ as a function of both the event horizon radius $r_{h}$ and the scale-dependent parameter $\gamma$, with fixed parameter values $\beta=2$ and $k=1$. The visualization clearly demonstrates that the energy increases monotonically with $r_{h}$, consistent with the thermodynamic interpretation of BH mass. The contour structure reveals that $\gamma$ exerts a negligible influence on $E$ at larger horizon scales, but becomes increasingly important in the small $r_{h}$ regime. This behavior indicates that scale-dependent geometric corrections subtly modify the BH energy structure near the Planckian regime, suggesting that quantum corrections encoded in $\gamma$ become critically relevant in the small-scale limit and may significantly influence evaporation dynamics or remnant formation scenarios.
Figure 1: Density plot of the inertial energy $E$ as a function of the event horizon radius $r_{h}$ and the scale-dependent parameter $\gamma$, with fixed values $\beta=2$ and $k=1$. The energy increases rapidly with $r_{h}$, while the influence of $\gamma$ is only significant in the small $r_{h}$ regime, highlighting the minimal but non-negligible role of scale-dependent corrections at short distances.
The Helmholtz free energy, which provides crucial information about the system’s equilibrium properties, is defined through the standard thermodynamic relation
[…]
yielding upon substitution and integration
$F\approx-\frac{\pi k\,r^{5}}{20}-\frac{3\pi\,\beta^{2}\gamma r^{2}}{8}+\frac{\pi\beta r^{2}}{4}-\frac{kr}{2\pi}+\frac{3\beta^{2}\gamma}{4\pi r^{2}}-\frac{\beta}{2\pi r^{2}}.$ (37)
Figure 2 displays the Helmholtz free energy behavior, revealing the complex interplay between geometric parameters and thermodynamic stability. The QC pressure, fundamental for understanding the mechanical properties of the BH system, follows from the thermodynamic relation [86]
$P=-\frac{dF}{dV},$ (38)
which evaluates to the expression
$P\approx\frac{r^{2}\pi k}{16\pi}+\frac{3\pi\,\beta^{2}\gamma}{16\pi r}-\frac{\pi\beta}{8\pi r}+\frac{k}{8\pi r^{2}\pi}+\frac{3\beta^{2}\gamma}{8\pi r^{5}\pi}-\frac{\beta}{4\pi r^{5}\pi}.$ (39)
Figure 2: Helmholtz free energy $F$ displaying the QC thermodynamic landscape for CWGBHs, showing the complex parameter dependence that governs equilibrium stability conditions.
Figure 3: QC pressure $P$ as a function of horizon parameters, illustrating the mechanical response properties of CWGBHs under varying geometric conditions.
The enthalpy, incorporating quantum corrections and representing the total energy content under constant pressure conditions, follows as
$H=E+PV,$ (40)
resulting in the detailed expression
$H\approx\frac{17\pi k\,r^{5}}{60}-\frac{\pi\,\beta^{2}\gamma r^{2}}{2}+\frac{\pi\gamma r^{4}}{8}+\frac{\pi\beta r^{2}}{3}\\ +\frac{kr}{6\pi}+\frac{\beta^{2}\gamma}{2\pi r^{2}}-\frac{\beta}{3\pi r^{2}}.$ (41)
Figure 4: Enthalpy $H$ distribution demonstrating the energy content of CWGBHs under constant pressure conditions, revealing the quantum-enhanced thermodynamic structure.
The Gibbs free energy, essential for assessing global thermodynamic stability and phase transition behavior, becomes
$G\approx\frac{\pi k\,r^{5}}{30}-\frac{\pi\,\beta^{2}\gamma r^{2}}{8}+\frac{\pi\beta r^{2}}{12}-\frac{kr}{3\pi}+\frac{5\beta^{2}\gamma}{4\pi r^{2}}-\frac{5\beta}{6\pi r^{2}}.$ (42)
Figure 5: Gibbs free energy $G$ characterizing the global thermodynamic stability landscape and phase transition structure of QC CWGBHs.
Finally, the QC heat capacity, which governs the thermal response and phase transition behavior, is expressed through the fundamental thermodynamic relation [49]
$C=T_{H}\left(\frac{\partial S}{\partial T_{H}}\right),$ (43)
leading to the final result
$C\approx\frac{\pi^{2}\left(2k\,r^{3}-3\beta^{2}\gamma+\gamma r^{2}+2\beta\right)r^{4}}{k\,r^{3}+3\beta^{2}\gamma-2\beta}.$ (44)
Figure 6 presents the heat capacity $C$ as a function of the event horizon radius $r_{h}$ and the scale-dependent parameter $\gamma$, revealing critical thermodynamic behavior that characterizes phase transitions in CWGBHs. The color gradient in this contour plot distinctly separates regions of negative and positive heat capacity, corresponding to thermodynamically unstable and stable phases, respectively. For low values of $\gamma$, a sharp transition from negative to positive $C$ occurs as $r_{h}$ increases, marked by a divergence in the heat capacity that signals a second-order phase transition. This behavior becomes less pronounced as $\gamma$ increases, demonstrating that scale-dependent corrections tend to smooth out the phase structure and can delay or suppress thermodynamic instabilities. The critical region near $r_{h}\approx 1.4$ and $\gamma\lesssim 0.2$ identifies a threshold horizon radius beyond which the BH becomes thermodynamically stable.
Figure 6: Heat capacity $C$ as a function of the event horizon radius $r_{h}$ and the scale-dependent parameter $\gamma$, with fixed parameters $\beta=2$ and $k=1$. The plot reveals a critical behavior in BH thermodynamics, where the sign change and divergence in $C$ for low $\gamma$ values indicate second-order phase transitions separating thermodynamically unstable (blue) and stable (red) regions.
VI QC JTE in CWGBHs
The JTE represents a fundamental thermodynamic process characterized by constant enthalpy conditions while allowing temperature variations in response to pressure changes. Within the framework of BH thermodynamics, particularly in the extended phase space formalism that treats the cosmological constant as a thermodynamic pressure, JTE provides a powerful tool for examining isenthalpic evolutionary pathways of gravitational systems [88, 35, 89]. The central diagnostic quantity for this analysis is the Joule-Thomson coefficient, which serves as a critical indicator determining whether a BH undergoes cooling or heating during pressure-driven expansion processes under constant enthalpy conditions.
The heat capacity, which plays an essential role in JTE analysis, has been systematically derived through application of the first law of BH thermodynamics using QC expressions for both entropy and temperature, as demonstrated in Eq.(43) and culminating in the comprehensive result given by Eq.(44). This QC heat capacity captures the subtle modifications introduced by CWG geometry and provides the foundation for understanding thermal response properties in the quantum gravitational regime [90].
The Joule-Thomson coefficient $\mu_{J}$ is rigorously defined as the partial derivative of temperature with respect to pressure under constant enthalpy conditions
$\mu_{J}=\left(\frac{\partial T_{H}}{\partial P_{C}}\right)_{H},$ (45)
where the subscript $H$ explicitly denotes the constraint of constant enthalpy. The physical interpretation of this coefficient is straightforward yet profound: positive values of $\mu_{J}$ correspond to cooling regimes where temperature decreases with increasing pressure, while negative values signal heating regimes where temperature increases under pressure enhancement.
This coefficient admits an alternative representation through fundamental thermodynamic relations, expressed as
$\mu_{J}=\frac{1}{C}\left[T_{H}\left(\frac{\partial V}{\partial T_{H}}\right)_{P}-V\right],$ (46)
where $C$ represents the QC heat capacity derived previously, and $V$ denotes the effective thermodynamic volume associated with the BH system. This formulation explicitly reveals the connection between thermal expansion properties and the Joule-Thomson effect, highlighting the role of heat capacity in mediating the thermodynamic response.
For practical computational purposes, particularly when dealing with the complex expressions arising in CWG, it proves convenient to express $\mu_{J}$ in terms of the horizon radius $r_{h}$ through the chain rule application
$\mu_{J}=\left(\frac{\partial T_{H}}{\partial r_{h}}\right)\cdot\left(\frac{\partial P}{\partial r_{h}}\right)^{-1}.$ (47)
Applying this computational framework to the QC thermodynamic quantities derived for CWGBHs, we obtain the explicit analytical expression
$\mu_{J}\approx\frac{8\pi r^{3}\left(kr^{3}+3\beta^{2}\gamma-2\beta\right)}{3\gamma\beta^{2}\left(-\pi^{2}r^{4}-10\right)+2\beta\left(\pi^{2}r^{4}+10\right)+2kr^{3}\left(\pi^{2}r^{4}-2\right)}$ (48)
This complex expression encapsulates the rich interplay between the various CWG parameters ($k$, $\gamma$, $\beta$) and the horizon geometry, revealing how quantum gravitational corrections modify the classical Joule-Thomson behavior. The explicit forms for the derivatives $\partial T_{H}/\partial r_{h}$ and $\partial P_{C}/\partial r_{h}$ emerge directly from the specific QC thermodynamic relations characteristic of the CWG model, yielding analytical or semi-analytical results depending on the computational complexity of the corrected expressions [91].
Comprehensive numerical analysis reveals the intricate behavior of $\mu_{J}$ as a function of the radius of the horizon $r_{h}$ in different parameter regimes, particularly focusing on variations in the scale-dependent parameter $\gamma$ and the cosmological parameter $k$. The transition between the cooling and heating regimes manifests itself as clear sign changes in $\mu_{J}$, providing an unambiguous identification of the boundaries of the thermodynamic phase. These transitions are not merely mathematical artifacts but represent genuine physical phenomena where the BH thermal response fundamentally changes character, capturing the QC phase structure inherent to CWGBHs under JTE conditions.
Figure 7 presents a comprehensive visualization of the JTE coefficient $\mu_{J}$ as a function of the event horizon radius $r_{h}$ for various values of the scale-dependent parameter $\gamma$, with fixed parameters $\beta=2$ and $k=1$. This plot elegantly illustrates the BH thermal response under isenthalpic expansion conditions, where the sign of $\mu_{J}$ serves as the crucial diagnostic: positive values indicate cooling phases during expansion, while negative values correspond to heating phases. For the baseline case $\gamma=0$, the transition from heating to cooling regimes occurs smoothly without dramatic features. However, as $\gamma$ increases, the curves develop pronounced vertical asymptotes that correspond to divergence points in $\mu_{J}$. These divergences mark critical horizon radii where the BH undergoes Joule-Thomson inversion, representing fundamental changes in the thermal response mechanism.
The systematic shift of these inversion points toward larger $r_{h}$ values with increasing $\gamma$ reveals a profound physical insight: the scale-dependent modifications characteristic of CWG effectively delay the onset of cooling behavior and introduce increasingly complex phase structure into the BH thermodynamics. This behavior demonstrates the remarkable sensitivity of BH thermal properties to quantum-inspired corrections and strongly supports the interpretation that the parameter $\gamma$ functions as a control parameter governing both inversion temperature and associated phase transition characteristics. The rich structure revealed in this analysis emphasizes how CWG provides a natural framework for understanding quantum gravitational effects in BH thermodynamics beyond the limitations of classical GR.
Figure 7: JTE coefficient $\mu_{J}$ versus event horizon radius $r_{h}$ for different values of the scale-dependent parameter $\gamma$, with fixed $\beta=2$ and $k=1$. The divergence points of $\mu_{J}$ indicate Joule-Thomson inversion radii, separating regions of heating ($\mu_{J}<0$) from cooling ($\mu_{J}>0$). As $\gamma$ increases, these inversion points shift to larger $r_{h}$, reflecting the impact of scale-dependent geometry on the thermodynamic response of the BH.
VII Gravitational Redshift in CWGBH Geometry
Gravitational redshift represents one of the most profound and experimentally verified predictions of GR, arising fundamentally from the curvature of spacetime in the vicinity of massive gravitational sources such as BHs. When electromagnetic radiation is emitted from regions characterized by strong gravitational fields, it undergoes a systematic shift toward longer wavelengths as it propagates away from the gravitational source, escaping to regions of weaker field strength [92, 93]. This phenomenon provides invaluable observational information about the underlying geometric structure of spacetime and serves as a powerful diagnostic tool for probing the nature of gravitational fields in both classical and quantum regimes.
In the context of CWG, where the metric structure deviates significantly from standard Schwarzschild geometry due to higher-order curvature corrections, the gravitational redshift acquires additional complexity and richness that reflects the underlying conformal symmetry structure [3]. The study of redshift in CWGBHs, therefore, offers unique insights into the observational signatures of alternative gravity theories and provides potential avenues for distinguishing CWG from conventional GR through precision astronomical observations.
For the analysis of light propagation and redshift phenomena, we focus on static, spherically symmetric configurations where angular variations can be safely neglected by setting $\theta$ and $\phi$ to constant values. Under these assumptions, the relevant portion of the metric line element simplifies to the two-dimensional form
$ds^{2}=B(r)dt^{2}-\frac{dr^{2}}{B(r)},$ (49)
where $B(r)$ represents the radial metric function derived from the CWG field equations, incorporating the combined gravitational influence of the BH mass parameter $\beta$, the scale-dependent correction $\gamma$, and the cosmological parameter $k$. This metric function encapsulates the essential departures from Schwarzschild geometry that characterize the CWG framework.
For massless particles following null geodesics, which represent the trajectories of photons propagating through the curved spacetime, the line element satisfies the fundamental constraint $ds^{2}=0$. Applying this condition to the simplified metric yields the geodesic equation
$\dot{r}=\pm\sqrt{B(r)},$ (50)
where $\dot{r}$ denotes the radial velocity component of light rays, with the sign determining whether the photon is propagating inward (negative) or outward (positive) relative to the gravitational source. This expression forms the fundamental foundation for evaluating redshift phenomena in the CWGBH geometry and provides the essential connection between the metric structure and observable electromagnetic effects.
The gravitational redshift parameter $z$, which quantifies the fractional change in photon frequency between emission and observation points, can be systematically linked to the underlying metric structure through the well-established relation [49]
$z=\left[1-2\dot{r}\right]^{1/2}-1.$ (51)
This formula encapsulates the fundamental physical principle that gravitational redshift arises from the difference in gravitational potential between emission and observation points, with the metric function serving as the mathematical representation of this potential difference in the curved spacetime geometry.
Substituting the explicit form of the CWGBH metric function into this expression and performing the necessary algebraic manipulations, we obtain the complete analytical expression for gravitational redshift in CWG geometry
$z=\sqrt{-1+6\beta\gamma+\frac{2\beta(2-3\beta\gamma)}{r}-2\gamma r-2kr^{2}}-1.$ (52)
This remarkable expression reveals the complex interplay between all CWG parameters in determining the redshift characteristics. The presence of terms involving $\beta\gamma$ coupling, linear $\gamma r$ contributions, and quadratic $kr^{2}$ dependencies demonstrates how the higher-order nature of CWG field equations manifests in observable electromagnetic phenomena. Unlike the simple $1/r$ dependence characteristic of Schwarzschild redshift, the CWG result exhibits a rich radial structure that could potentially be detected through high-precision spectroscopic observations of radiation emanating from the vicinity of massive compact objects.
Figure 8 presents a comprehensive visualization of the gravitational redshift behavior in CWGBH geometry, illustrating how the redshift parameter $z$ varies as a function of radial distance and the various CWG parameters. The plot reveals several distinctive features that distinguish CWG redshift from its GR counterpart: the presence of oscillatory behavior at intermediate distances due to the linear $\gamma r$ term, the modified asymptotic behavior at large distances governed by the cosmological parameter $k$, and the enhanced redshift effects near the horizon arising from the complex parameter couplings. These features provide potentially observable signatures that could serve as discriminating tests between CWG and standard GR, offering new avenues for testing alternative gravity theories through electromagnetic observations of compact astrophysical objects such as neutron stars, stellar-mass BHs, and supermassive BHs in galactic centers.
Figure 8: Gravitational redshift $z$ in CWGBH geometry showing the complex radial dependence arising from CWG corrections. The plot demonstrates how the redshift behavior deviates from standard Schwarzschild predictions due to the presence of linear $\gamma r$ and quadratic $kr^{2}$ terms, providing potentially observable signatures for distinguishing CWG from conventional GR through precision spectroscopic measurements.
VIII Conclusion
In this comprehensive theoretical investigation, we systematically explored the thermodynamic and quantum features of BHs arising from the Mannheim-Kazanas solution within the framework of CWG. Unlike the conventional Einstein-Hilbert paradigm based on second-order field equations, CWG employs a fourth-order action with local conformal symmetry, leading to fundamentally modified gravitational dynamics and significantly richer vacuum structures [66, 94]. The inclusion of additional potential terms, specifically the linear $\gamma r$ and quadratic $kr^{2}$ contributions in the metric function given by Eq.(10), enabled an alternative theoretical description of gravitational phenomena spanning both galactic and cosmological scales without requiring the introduction of dark matter components.
Our analysis commenced with the systematic application of the Hamilton-Jacobi tunneling formalism to compute the Hawking temperature of CWGBHs, as detailed in Section III. This investigation revealed explicit contributions from all conformal parameters $\beta$, $\gamma$, and $k$, culminating in the comprehensive temperature expression presented in Eq.(26). The resulting thermal spectrum deviates markedly from that of standard Schwarzschild-AdS BHs, particularly due to the influence of the linear potential term $\gamma r$, demonstrating the remarkable sensitivity of BH radiation to conformal corrections [95, 96]. Table 1 provided quantitative evidence of these effects, showing temperature enhancements exceeding 75,000% in certain parameter regimes compared to baseline Schwarzschild values.
We subsequently incorporated quantum gravitational effects through the GUP framework, deriving the QC Hawking temperature expression that accounts for minimal length scale effects. This analysis demonstrated that the presence of a fundamental length scale systematically suppresses the thermal spectrum, particularly in the near-Planckian regime where $r_{h}\to l_{p}$. The resulting temperature suppression reflects a gravitational redshift in the emitted radiation and suggests the potential emergence of remnant-like structures during the final stages of BH evaporation [97, 98]. This quantum modification provides crucial insights into the transition between semiclassical and fully quantum gravitational regimes.
In Section V, we employed an exponentially corrected entropy model, as expressed in Eq.(33), to systematically compute the complete set of QC thermodynamic potentials for CWGBHs. These calculations yielded detailed expressions for internal energy (Eq.(35)), Helmholtz free energy (Eq.(37)), pressure (Eq.(39)), enthalpy (Eq.(41)), and heat capacity (Eq.(44)). Figure 1 illustrates the complex energy landscape, revealing how scale-dependent corrections become increasingly important in the small horizon radius regime. Most significantly, Figure 6 demonstrated the existence of thermodynamic phase transitions through the heat capacity analysis, showing distinct regions of negative and positive heat capacity that correspond to thermodynamically unstable and stable phases, respectively.
The JTE analysis presented in Section VI provided a sophisticated probe of isenthalpic processes in CWGBH thermodynamics within the extended phase space formalism. The behavior of the Joule-Thomson coefficient $\mu_{J}$ revealed distinct cooling and heating phases that depend sensitively on both the horizon radius and the CWG parameters. Figure 7 captured the QC phase structure, showing how the scale-dependent parameter $\gamma$ controls the location of Joule-Thomson inversion points and introduces increasingly complex thermal response behavior. These results demonstrated that CWG naturally accommodates rich phase structures that are absent in conventional GR treatments.
Our investigation of gravitational redshift within CWGBH geometry, detailed in Section VII, revealed that the frequency shift of electromagnetic radiation emitted near these objects encodes valuable information about the underlying conformal metric structure. The derived redshift expression showed complex radial dependence arising from the interplay of all CWG parameters, providing potentially observable signatures that could serve as discriminating tests between CWG and standard GR through precision spectroscopic observations. Figure 8 shows how the redshift behavior deviates systematically from Schwarzschild predictions, offering new avenues for experimental tests of alternative gravity theories.
Throughout our analysis, we demonstrated that CWG offers a remarkably consistent and theoretically rich framework for studying BH thermodynamics beyond the limitations of GR, particularly when augmented with quantum corrections that become essential near the Planck scale. The intricate interplay between higher-derivative gravitational dynamics and quantum-statistical effects opened promising new avenues toward understanding quantum gravity phenomenology in astrophysically relevant contexts [99]. Our results provide both theoretical foundations and practical computational tools for investigating quantum gravitational effects in realistic BH systems.
Looking toward future research directions, our comprehensive theoretical framework suggests several promising avenues for extended investigation. First, the development of more sophisticated numerical techniques for solving the complete nonlinear CWG field equations would enable precise predictions of observational signatures, including gravitational wave emission patterns, accretion disk spectra, and electromagnetic radiation characteristics that could be compared with data from current and future astronomical facilities. Second, the extension of our thermodynamic analysis to rotating CWGBHs would provide crucial insights into the effects of angular momentum on QC thermal properties, potentially revealing new classes of phase transitions and instabilities relevant to realistic astrophysical scenarios. Third, comprehensive observational studies involving gravitational lensing phenomena, particularly strong lensing effects around supermassive BHs, could provide direct constraints on the CWG parameters $\beta$, $\gamma$, and $k$ through comparison with high-resolution imaging data from facilities such as the Event Horizon Telescope. Fourth, the investigation of gravitational wave signatures from CWGBHs during merger events could yield distinctive waveform characteristics that differ from GR predictions, offering new opportunities for testing alternative gravity theories through advanced LIGO/Virgo observations and future space-based detectors.
Finally, another study of cosmological implications of CWG, particularly its role in explaining dark matter phenomena and cosmic acceleration without exotic matter components, might represent an extension of our BH-focused investigation. Therefore, our future research agenda is set to enhance our grasp of quantum gravity phenomena and offer empirical methods for examining the essential characteristics of spacetime geometry within extreme astrophysical environments.
Acknowledgements.
We express gratitude to Dr. Ercan Kılıçarslan and Dr. Heeseung Zoe for their constructive feedback. İ. S. appreciates the academic support from EMU, TÜBİTAK, ANKOS, and SCOAP3, as well as the networking opportunities through COST Actions CA22113, CA21106, and CA23130.